INFRARED SPECTRA OF NEUTRAL AND IONIC SO_{2}H_{2} SPECIES TRAPPED IN SOLID NEON

\underline{MARILYN E. JACOX} AND WARREN E. THOMPSON, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

When a Ne:H_{2}:SO_{2} = 600:10:1 mixture was codeposited at 4.2 K with a beam of neon atoms that had been excited in a microwave discharge, the infrared spectrum of the resulting solid included new absorptions contributed by isomers of S(OH)_{2} and of its cation. In other experiments in which the neon atoms were not excited in a discharge but in which the SO_{2} was excited by 254 nm radiation, isomers of uncharged S(OH)_{2} and of HS(O)OH appeared. Isotopic substitution studies and density functional calculations support the infrared identifications.